This is not a financial question however, it deals with percentages. So I presume this is the best board to post to. I apologize in advance if it is not.
Soybean meal is 12% protein; cornmeal is 6% protein. How many pounds of each should be mixed together in order to get 240-lb mixture that is 8% protein?
How many pounds of cornmeal should be in the mixture?
___ pounds
How many pounds of soybean meal should be in the mixture?
___ pounds
I dont understand this at all. The protein percentages do not have a value for how many pounds (of each) it takes to come up with the 12% and 6%

Answer by stanbon(75887)

Soybean meal is 12% protein; cornmeal is 6% protein. How many pounds of each should be mixed together in order to get 240-lb mixture that is 8% protein?
-------
Using 2 variables.
Equation:
Quantity: s + c = 240
Protein:0.12s+0.06c = 0.08*240
-----
Multiply thru 1st by 12
Multiply thru 2nd by 100
-------------------
12s + 12c = 12*240
12s + 6c = 8*240
-----------------
Subtract and solve for "c":
6c = 4*240
c = (2/3)(240)
c = 160 lb (amt. of cornmeal needed)
---
Solve for "s":
s+c = 240
s+160 = 240
s = 80 lb (amt of soybean needed)
=======================================
Cheers,
Stan H.
===========



How many pounds of cornmeal should be in the mixture?
___ pounds
How many pounds of soybean meal should be in the mixture?
___ pounds